# Evolving the Euler rotation axis as a dynamical system, using the Euler vector and generalizations

### v1

@@ Abstract.
Euler’s theorem, rigid body motion, axis-angle, spinors, quasiperiodicity
@@
Existence of angular velocity. For any rigid body motion that leaves the origin of the body system fixed, there exists a unique (axial) vector function $\boldsymbol{\omega}(t)$ such that for any point in the body, if $\bf{r}(t)$ is its displacement at time $t$, then
@@
We often suppress $t$ and write ${\bf{\dot{r}}}=\boldsymbol{\omega}\times\bf{r}$. If the points of the body have continuous velocities, $\boldsymbol{\omega}$ is continuous.
The existence of angular velocity is either assumed or proved (with varying amounts of care) in every physics book (e.g. [2] p. 172). A rigorous proof shows that $\boldsymbol{\omega}(t)$ is continuous if the points in the body have continuous velocities. We shall assume $\boldsymbol{\omega}(t)$ is continuous.
Let $d\boldsymbol{\theta}=\boldsymbol{\omega}dt$, so $d{\bf{r}}=d\boldsymbol{\theta}\times{\bf{r}}$. Thus, infinitesimal motion is indeed rotation, with rotation axis determined by the direction of $d\boldsymbol{\theta}$, and angle $\left|{d{\boldsymbol{\theta}}}\right|$ counterclockwise using the right-hand rule, with thumb pointing in direction of $\boldsymbol{\omega}$. Everyone is happy with calling this a vector, more precisely an axial vector, representing the infinitesimal motion.
For non-infinitesimal rotations of a rigid body, one can do the same thing. Suppose ${\boldsymbol{\hat{E}}}$ is a unit vector pointing along the axis of the rotation, and $\theta$ is the counterclockwise angle of rotation about that axis. Define a finite analog of the infinitesimal rotation vector as
@@
This is called an Euler vector in ([14], [15]) and we will use that terminology. It is also called a rotation vector ([16], [17],[14]), or axis-angle vector [18]. It is referred to as the Rodrigues parameter in [16], though that term is defined differently in [18]. But one thing they all do is apologize for calling this a vector. If $\boldsymbol{E}_{1}$ and $\boldsymbol{E}_{2}$ are Euler vector representations of two rotations, then $\boldsymbol{E}_{1}+\boldsymbol{E}_{2}$ is generally not an Euler vector representation of the result of doing the two rotations in succession (in either order). On account of this, [16] p. 236, says that the Euler vector is not a “real vector”, and [14] says this “shows that finite rotations are not really vectors at all”. The physics books don’t even write it down; [2], p. 163, says that “such a correspondence cannot be made successfully”.
@@
We disagree. While it is true that finite rotations “are” not literally vectors (they are not literally matrices either: they are operations), that doesn’t mean they can’t be represented by vectors. One property of the Euler vector that attests to its vector quality is that when drawn as an arrow in the real physical space of the rigid body, one sees that it is more than some arbitrary parameterization of the three degrees of freedom of the body: one can view the rotation being around the axis defined by that arrow in the picture, so its direction has real meaning, not an abstraction. Another test of being a vector is how it transforms. Under an orthogonal transformation, the rotation axis for a rigid body clearly just transforms along with the body. If the orthogonal transformation is orientation-preserving, the counterclockwise sense of rotation stays the same, so the Euler vector is transformed the same way as the vectors in the body. If the transformation is not orientation-preserving (it’s determinant is -1), the transformed Euler vector has to be flipped, because the meaning of counterclockwise got reversed. In other words, the Euler vector transforms as an axial vector, just like angular velocity does.
@@
Finally, an analogy: the unit vectors are a subset of $\mathbb{R}^{3}$ (the unit ball). The sum of two of them is not a unit vector, but no one has ever said that unit vectors aren’t "really" vectors on account of that. They form a subset, not a subspace, but they are still vectors, elements of $\mathbb{R}^{3}$ But if the reader prefers to insist that the Euler vector is not truly a vector, it will not affect the discussion that follows. The reason we bring it up is that perhaps the dynamics of the Euler vector had not been previously considered due to its vector status being discounted. But it is a natural quantity for passing from infinitesimal to finite motions, because the vector that is always used to represent an infinitesimal motion is already in Euler vector form. In the next section we also consider generalizations of the Euler vector which behave just like the Euler vector in the infinitesimal limit.
@@
Let $Rot_{{\bf{n}},\theta}$ be the operation counterclockwise rotation by angle $\theta$ around the axis through the origin with direction unit vector $\bf{n}$; if $\theta$ is a multiple of $2\pi$, it is understood to be the identity operator even if $\bf{n}$ is undefined. The rotation can be represented by Euler vector $\boldsymbol{E}=\bf{n}\theta$. In ([14], [15]), $\theta$ is called the length of the vector, which would make it non-negative. But $Rot_{{\bf{n}},\theta}$ makes sense when $\theta$ is negative; a counterclockwise rotation by a negative amount is just a clockwise rotation by that amount. Since ${\bf{n}}\theta=({-\bf{n}})(-\theta)$, $\boldsymbol{E}=\bf{n}\theta$ could always be written as a unit vector times the length of $\bf{E}$, with the length representing the rotation angle, if desired. But later we’ll have occasion to need $\theta$ to be negative, in describing a continuous evolution of the Euler rotation axis and angle.
@@
Rodrigues’ rotation formula. For any point $\bf{r}$,
@@
To find the total $d\boldsymbol{F}$, since the motions are infinitesimal, we can add together the cases for ${\boldsymbol{\omega}}$ parallel and perpendicular to $\boldsymbol{F}$. In general, write ${\boldsymbol{\omega}}={{\boldsymbol{\omega}}_{\parallel}}+{{\boldsymbol{\omega}}_{\bot}}={\bf{n}}\left({{\boldsymbol{\omega}}\cdot{\bf{n}}}\right)+\left({{\boldsymbol{\omega}}-{\bf{n}}\left({{\boldsymbol{\omega}}\cdot{\bf{n}}}\right)}\right)$. Writing it as a differential equation, $\boldsymbol{\dot{F}}=d\boldsymbol{F}/dt={\bf{n}}\left({{\boldsymbol{\omega}}\cdot{\bf{n}}}\right)f^{\prime}(\theta)+\left({{\boldsymbol{\omega}}-{\bf{n}}\left({{\boldsymbol{\omega}}\cdot{\bf{n}}}\right)}\right)f(\theta)\cot(\theta/2)/2+{\boldsymbol{\omega}}\times\boldsymbol{F}/2$. Rearranging terms, we have:
Generalized Euler vector equation.
@@
Euler axis and angle as separate variables, and a pathological example. Consider the behavior of $\bf{n}$ and $\theta$. Suppose ${\bf{F}}(t)$ satisfies (4) on some interval $t_{0}\leq t<t_{1}$, and that $0<|{\bf{F}}(t)|\leq K$ for $t_{0}<t<t_{1}$, so that ${\bf{F}}(t)={\bf{n}}(t)f(\theta(t))$ uniquely determines ${\bf{n}}(t)$ and $\theta(t)>0$ as continuous functions on $t_{0}<t<t_{1}$. (Or, the sign of $\theta(t)$ and ${\bf{n}}(t)$ could both be flipped, but we don’t need to discuss negative $\theta$ until a later section). If $|{\bf{F}}(t_{0})|=0$, then $\theta(t_{0})=0$ and ${\bf{n}}(t_{0})$ is undefined. Now $f^{2}(\theta)={\bf{F}}\cdot{\bf{F}}$, so $2f(\theta)f^{\prime}(\theta){\dot{\theta}}=2{\bf{F}}\cdot\dot{{\bf{F}}}$, and from (4), ${\bf{F}}\cdot\dot{{\bf{F}}}={\boldsymbol{\omega}}\cdot{\bf{F}}f^{\prime}(\theta)$. Neither $f(\theta)$ or $f^{\prime}(\theta)$ is zero, so $\dot{\theta}={\boldsymbol{\omega}}\cdot{\bf{n}}$. Also, differentiating ${\bf{F}}={\bf{n}}f(\theta)$ gives ${\dot{\bf{F}}}={\dot{\bf{n}}}f(\theta)+{\bf{n}}f^{\prime}(\theta)\dot{\theta}$. This may be solved for ${\dot{\bf{n}}}$, using (4) and $\dot{\theta}={\boldsymbol{\omega}}\cdot{\bf{n}}$. For future reference, we record these, for $t_{0}<t<t_{1}$:
@@
The fact that $\boldsymbol{F}(t)$ behaves nicely as $t\rightarrow 0^{+}$ for any continuous $\boldsymbol{\omega}$, when $|{\bf{F}}(0)|=0$, does not imply that ${\bf{n}}(t)$ does. Here is a pathological example. Let
@@
Euler vector equation. Take $f(\theta)=\theta$, and refer to $\bf{F}$ as $\bf{E}$. Then the differential equation, written entirely in terms of $\bf{E}$, becomes
@@
Modified Gibbs vector equation, and quaternions. A glance at (4) shows that the equation is by far the simplest if we take $f(\theta)=\sin(\theta/2)$, because the middle term disappears entirely. The vector ${\bf{M}}={\bf{n}}\sin(\theta/2)$ is called the modified Gibbs representation in [16], motivated in a different way. The differential equation becomes
@@
Gibbs vector equation. The representation ${\bf{G}}={\bf{n}}\tan(\theta/2)$ is called called variously the Gibbs vector [16], or the Rodrigues vector [14]. Differential equation (4) becomes
@@
Vector identity ([20] p. 77, also [21]). For ${\bf{A,B,C,D}}$ in $\mathbb{R}^{3}$,
@@
Theorem 1 accomplishes what initially motivated us: to use only infinitesimal motions to generate all future positions of the body, by means of a differential equation, and in the process showing that finite motions are also rotations, giving yet another proof of Euler’s rotation theorem. The version with ${\bf{M}}$ gives an axis and angle for the position for all time, though they do not necessarily evolve continuously for all time; example (6) shows that such is not always possible. The version with ${\bf{E}}$ also could be used to verify all future positions are rotations, even though the solution for ${\bf{E}}$ cannot necessarily be continued for all time, if $|{\bf{E}}(t)|$ approaches a positive multiple of $2\pi$ in finite time. If $|{\bf{E}}(t)|$ reaches say $23\pi/$ 2 at time $t_{1}$, restart with initial condition $-{\bf{E}}(t_{1})/3$, which amounts to flipping the axis and starting over with rotation angle $\pi/2$, corresponding to the same physical position. This avoids dealing with the boundary issues.
@@
$=1-\left({1-{{({\boldsymbol{\omega}}\cdot{{\bf{n}}_{0}})}^{2}}}\right){\sin^{2}}({\theta_{0}}/2)={a^{2}},$ so $\left|{{{\bf{v}}_{2}}}\right|=1,$ and so ${{\bf{v}}_{2}}={{\bf{n}}_{0}}({\boldsymbol{\omega}}\cdot{{\bf{n}}_{0}}/a)+{\bf{u}}\sqrt{1-{{({\boldsymbol{\omega}}\cdot{{\bf{n}}_{0}}/a)}^{2}}}.$ Similarly, ${{\bf{n}}_{0}}\sqrt{1-{{({\boldsymbol{\omega}}\cdot{{\bf{n}}_{0}})}^{2}}}\cos({\theta_{0}}/2)-{\bf{u}}({\boldsymbol{\omega}}\cdot{{\bf{n}}_{0}})$ has length $a$, so ${\left({1-{a^{2}}}\right)^{-1/2}}{{\bf{v}}_{1}}={{\bf{n}}_{0}}\sqrt{1-{{({\boldsymbol{\omega}}\cdot{{\bf{n}}_{0}}/a)}^{2}}}-{\bf{u}}({\boldsymbol{\omega}}\cdot{{\bf{n}}_{0}}/a).$ Thus ${\bf{n}}(t)\sin({\theta(t)}/2){\left({1-{a^{2}}}\right)^{-1/2}}$$={{\bf{e}}_{1}}\cos(t/2+b)+{{\bf{e}}_{2}}{\left({1-{a^{2}}}\right)^{-1/2}}\sin(t/2+b)={\bf{p}}(t),$ and ${\bf{n}}(t)$ is the normalization of this. ∎
Example. Suppose the body starts out from having been rotated 90 degrees about the $x$ axis, and thereafter has angular velocity vector ${\bf{k}}$, so it is spinning about the vertical. That is, ${\theta_{0}}=\pi/2$, ${{\bf{n}}_{0}}={\bf{i}}$, and ${\boldsymbol{\omega}}={\bf{k}}$, and let ${t_{0}}=0$. Then in Theorem 2, $a=1/\sqrt{2}$ , $b=0$, and $\theta(t)=2{{\mathop{\rm Cos}\nolimits}^{-1}}\left({{2^{-1/2}}\cos\left({t/2}\right)}\right)$. Also, one can compute ${{\bf{e}}_{1}}={\bf{i}}$ and ${{\bf{e}}_{2}}=({\bf{j}}+{\bf{k}})/\sqrt{2}$, so ${\bf{n}}(t)=(\cos(t/2),\sin(t/2),\sin(t/2))/\sqrt{1+{{\sin}^{2}}(t/2)}.$ This is illustrated above in Figures 1 and 2. Note ${\bf{n}}$ starts at ${{\bf{i}}}$, picks up some positive 2nd and 3rd components, and then winds up at $-{{\bf{i}}}$ as $\theta$ goes from $\pi/2$ to $23\pi/$ 2, as the body makes one complete revolution. Then ${\bf{n}}$ picks up some negative 2nd and 3rd components, and winds up at ${{\bf{i}}}$ as $\theta$ goes from $23\pi/$ 2 back to $\pi/2$, as the body makes another complete revolution. Any example of Theorem 2 behaves similarly, except that the offset $b$ is not zero unless ${\boldsymbol{\omega}}$ is perpendicular to ${{\bf{n}}_{0}}$, so in general $\bf{n}$ does not start out at ${{\bf{e}}_{1}}$. See Figure 5.
The spinor-like behavior of Theorem 2, in which ${\bf{n}}$ changes sign when the body makes one complete revolution, does not actually require that ${\boldsymbol{\omega}}$ be constant in magnitude, only in direction. This is consequence of a simple change of time argument that is analogous to changing to arclength parameter for a curve.
Lemma 3.
@@
Thus in Theorem 2, the hypothesis could be changed to ${\boldsymbol{\omega}}/\left|{\boldsymbol{\omega}}\right|$ being constant, and the conclusions would hold using the new time variable, and the statement about ${\bf{n}}$ changing sign every time the body makes one complete rotation is still true. In general, the geometrical behavior of $\boldsymbol{E}$ is not affected by a change of time scale, in the same way that the geometry of a curve is not affected by a change of parameterization. This is useful in our computer experiments with $\boldsymbol{E}$: the magnitude of ${\boldsymbol{\omega}}$ might as well be kept fixed.
@@
Theorem 4 (Continuous evolution of Euler axis and angle).
@@
We’ll prove the theorem by analyzing the solution to equation (9), which exists for all time, at points where ${\bf{M}}={\bf{0}}$. Under the hypotheses of Theorem 4, these points are isolated. The trick will be to choose the correct branch of the $\arccos$ function in Theorem 1 between these points so that axis and angle connect up continuously.
Lemma 5 (Zeros are isolated).
Let ${\boldsymbol{\omega}}$ satisfy the hypotheses of Theorem 4, and let $({M_{0}(t),\bf{M}}(t))$ solve equation (9) for ${t_{0}}\leq t$, with $M_{0}(t_{0})^{2}+{\bf{M}}(t_{0})\cdot{\bf{M}}(t_{0})=1$. If $t_{0}<t_{1}$ and ${\bf{M}}(t_{1})={\bf{0}}$, there exists $\delta>0$ such that ${\bf{M}}(t)\neq{\bf{0}}$ for $0<|t-t_{1}|<\delta$.
@@
Lemma 6 (Continuation at a zero).
Let ${\boldsymbol{\omega}}$ satisfy the hypotheses of Theorem 4, and let $({M_{0}(t),\bf{M}}(t))$ solve equation (9) for ${t_{0}}\leq t$, with $M_{0}(t_{0})^{2}+{\bf{M}}(t_{0})\cdot{\bf{M}}(t_{0})=1$. Suppose $t_{0}<t_{1}<t_{2}$, ${\bf{M}}(t_{1})={\bf{0}}$, and ${\bf{M}}(t)\neq{\bf{0}}$ if $t\neq t_{1}$ and $t_{0}<t<t_{2}$. Suppose, for some integer $l$, $\theta(t)$ has been defined for $t_{0}\leq t\leq t_{1}$ by $\cos(\theta(t)/2)=M_{0}(t)$ with $2\pi l\leq\theta(t)\leq 2\pi(l+1)$; and for $t_{0}<t<t_{1}$, ${\bf{n(t)}}={\bf{M}}(t)/\sin(\theta(t)/2)$. Let $i=min\{j\geq 0:{\boldsymbol{\omega}}^{(j)}(t_{1})\neq{\bf{0}}\}$.

### v2

@@ Abstract.
Euler’s rotation theorem, rigid body dynamics, axis-angle, spinors, quaternions, quasiperiodicity
@@
Lemma 2.1 (Existence of angular velocity).
For any rigid body motion that leaves the origin of the body system fixed, there exists a unique (axial) vector function $\boldsymbol{\omega}(t)$ such that for any point in the body, if $\bf{r}(t)$ is its displacement at time $t$, then
@@
We often suppress $t$ and write ${\bf{\dot{r}}}=\boldsymbol{\omega}\times\bf{r}$.
The existence of angular velocity is either assumed or proved (with varying amounts of care) in every physics book (e.g. [2] p. 172). A rigorous proof shows that $\boldsymbol{\omega}(t)$ is continuous if the points in the body have continuous velocities. We shall assume that $\boldsymbol{\omega}(t)$ is continuous.
Let $d\boldsymbol{\theta}=\boldsymbol{\omega}dt$, so $d{\bf{r}}=d\boldsymbol{\theta}\times{\bf{r}}$. Thus, infinitesimal motion is indeed rotation, with rotation axis determined by the direction of $d\boldsymbol{\theta}$, and angle $\left|{d{\boldsymbol{\theta}}}\right|$ counterclockwise using the right-hand rule, with thumb pointing in direction of $\boldsymbol{\omega}$. Everyone is happy to call this a vector, more precisely an axial vector, representing the infinitesimal motion.
For non-infinitesimal rotations of a rigid body, one can do the same thing. Suppose ${\boldsymbol{\hat{E}}}$ is a unit vector pointing along the axis of rotation, and $\theta$ is the counterclockwise angle of rotation about that axis. Define a finite analog of the infinitesimal rotation vector as
@@
This is called an Euler vector in ([14], [15]) and we will use that terminology. It is also called a rotation vector ([16], [17],[14]), or axis-angle vector [18]. It is referred to as the Rodrigues parameter in [16], although that term is defined differently in [18]. But one thing they all do is apologize for calling this a vector. If $\boldsymbol{E}_{1}$ and $\boldsymbol{E}_{2}$ are Euler vector representations of two rotations, then $\boldsymbol{E}_{1}+\boldsymbol{E}_{2}$ is generally not an Euler vector representation of the result of doing the two rotations in succession (in either order). On account of this, [16] p. 236, says that the Euler vector is not a "real vector", and [14] says that this "shows that finite rotations are not really vectors at all". The physics books don’t even write it down; [2], p. 163, says that “such a correspondence cannot be made successfully”.
@@
We disagree. Although it is true that finite rotations “are” not literally vectors (they are not literally matrices either: they are operations), that doesn’t mean they can’t be represented by vectors. One property of the Euler vector that attests to its vector quality is that when drawn as an arrow in the real physical space of the rigid body, one sees that it is more than some arbitrary parameterization of the three degrees of freedom of the body: one can view the rotation being around the axis defined by that arrow in the picture, so its direction has real meaning, not an abstraction. Another test of being a vector is how it transforms. Under an orthogonal transformation, the rotation axis for a rigid body clearly just transforms along with the body. If the orthogonal transformation is orientation-preserving, the counterclockwise sense of rotation stays the same, so the Euler vector is transformed the same way as the vectors in the body. If the transformation is not orientation-preserving (it’s determinant is -1), the transformed Euler vector has to be flipped, because the meaning of counterclockwise got reversed. In other words, the Euler vector transforms as an axial vector, just as does the angular velocity.
@@
Finally, an analogy: the unit vectors are a subset of $\mathbb{R}^{3}$ (the unit ball). The sum of two of them is not a unit vector, but no one has ever said that unit vectors aren’t "really" vectors on account of that. They form a subset, not a subspace, but they are still vectors, elements of $\mathbb{R}^{3}$ But if the reader prefers to insist that the Euler vector is not truly a vector, it will not affect the discussion that follows. The reason we bring it up is that perhaps the dynamics of the Euler vector had not been previously considered due to its vector status being discounted. But it is a natural quantity for passing from infinitesimal to finite motions because the vector that is always used to represent an infinitesimal motion is already in Euler vector form. In the next section, we also consider generalizations of the Euler vector which behave just like the Euler vector in the infinitesimal limit.
@@
Let $Rot_{{\bf{n}},\theta}$ be the operation of counterclockwise rotation by angle $\theta$ around the axis through the origin with direction unit vector $\bf{n}$; if $\theta$ is a multiple of $2\pi$, it is understood to be the identity operator even if $\bf{n}$ is undefined. The rotation can be represented by Euler vector $\boldsymbol{E}=\bf{n}\theta$. In ([14], [15]), $\theta$ is called the length of the vector, which would make it non-negative. But $Rot_{{\bf{n}},\theta}$ makes sense when $\theta$ is negative; a counterclockwise rotation by a negative amount is just a clockwise rotation by that amount. Since ${\bf{n}}\theta=({-\bf{n}})(-\theta)$, $\boldsymbol{E}=\bf{n}\theta$ could always be written as a unit vector times the length of $\bf{E}$, with the length representing the rotation angle, if desired. But later we’ll have occasion to need $\theta$ to be negative, in describing a continuous evolution of the Euler rotation axis and angle.
@@
Lemma 3.1 (Rodrigues’ rotation formula).
For any point $\bf{r}$,
@@
To find the total $d\boldsymbol{F}$, since the motions are infinitesimal, we can add together the cases for ${\boldsymbol{\omega}}$ parallel and perpendicular to $\boldsymbol{F}$. In general, write ${\boldsymbol{\omega}}={{\boldsymbol{\omega}}_{\parallel}}+{{\boldsymbol{\omega}}_{\bot}}={\bf{n}}\left({{\boldsymbol{\omega}}\cdot{\bf{n}}}\right)+\left({{\boldsymbol{\omega}}-{\bf{n}}\left({{\boldsymbol{\omega}}\cdot{\bf{n}}}\right)}\right)$. Writing it as a differential equation, $\boldsymbol{\dot{F}}=d\boldsymbol{F}/dt={\bf{n}}\left({{\boldsymbol{\omega}}\cdot{\bf{n}}}\right)f^{\prime}(\theta)+\left({{\boldsymbol{\omega}}-{\bf{n}}\left({{\boldsymbol{\omega}}\cdot{\bf{n}}}\right)}\right)f(\theta)\cot(\theta/2)/2+{\boldsymbol{\omega}}\times\boldsymbol{F}/2$. Rearranging terms, we find our first form of an Euler vector differential equation.
3.1. Generalized Euler vector equation.
@@
3.2. Euler axis and angle as separate variables.
Consider the behavior of $\bf{n}$ and $\theta$. Suppose ${\bf{F}}(t)$ satisfies (4) on some interval $t_{0}\leq t<t_{1}$, and that $0<|{\bf{F}}(t)|\leq K$ for $t_{0}<t<t_{1}$, so that ${\bf{F}}(t)={\bf{n}}(t)f(\theta(t))$ uniquely determines ${\bf{n}}(t)$ and $\theta(t)>0$ as continuous functions on $t_{0}<t<t_{1}$. (Or, the sign of $\theta(t)$ and ${\bf{n}}(t)$ could both be flipped, but we don’t need to discuss negative $\theta$ until a later section). If $|{\bf{F}}(t_{0})|=0$, then $\theta(t_{0})=0$ and ${\bf{n}}(t_{0})$ is undefined. Now $f^{2}(\theta)={\bf{F}}\cdot{\bf{F}}$, so $2f(\theta)f^{\prime}(\theta){\dot{\theta}}=2{\bf{F}}\cdot\dot{{\bf{F}}}$, and from (4), ${\bf{F}}\cdot\dot{{\bf{F}}}={\boldsymbol{\omega}}\cdot{\bf{F}}f^{\prime}(\theta)$. Neither $f(\theta)$ or $f^{\prime}(\theta)$ is zero, so $\dot{\theta}={\boldsymbol{\omega}}\cdot{\bf{n}}$. Also, differentiating ${\bf{F}}={\bf{n}}f(\theta)$ gives ${\dot{\bf{F}}}={\dot{\bf{n}}}f(\theta)+{\bf{n}}f^{\prime}(\theta)\dot{\theta}$. This may be solved for ${\dot{\bf{n}}}$, using (4) and $\dot{\theta}={\boldsymbol{\omega}}\cdot{\bf{n}}$. For future reference, we record these, for $t_{0}<t<t_{1}$:
@@
The fact that $\boldsymbol{F}(t)$ behaves nicely as $t\rightarrow 0^{+}$ for any continuous $\boldsymbol{\omega}$, when $|{\bf{F}}(0)|=0$, does not imply that ${\bf{n}}(t)$ does. Here is a pathological example.
Example.
Let
@@
3.3. Euler vector equation.
Take $f(\theta)=\theta$, and refer to $\bf{F}$ as $\bf{E}$. Then the differential equation, written entirely in terms of $\bf{E}$, becomes
@@
3.4. Modified Gibbs vector equation, and quaternions.
A glance at (4) shows that the equation is by far the simplest if we take $f(\theta)=\sin(\theta/2)$, because the middle term disappears entirely. The vector ${\bf{M}}={\bf{n}}\sin(\theta/2)$ is called the modified Gibbs representation in [16], motivated in a different way. The differential equation becomes
@@
3.5. Gibbs vector equation.
The representation ${\bf{G}}={\bf{n}}\tan(\theta/2)$ is called variously the Gibbs vector [16], or the Rodrigues vector [14]. Differential equation (4) becomes
@@
Lemma 3.3 (Vector identity).
([20] p. 77, also [21]). For ${\bf{A,B,C,D}}$ in $\mathbb{R}^{3}$,
@@
Theorem 3.2 accomplishes what initially motivated us: to use only infinitesimal motions to generate all future positions of the body, by means of a differential equation, and in the process showing that finite motions are also rotations, giving yet another proof of Euler’s rotation theorem. The version with ${\bf{M}}$ gives an axis and angle for the position for all time, though they do not necessarily evolve continuously for all time; example (6) shows that such is not always possible. The version with ${\bf{E}}$ also could be used to verify all future positions are rotations, even though the solution for ${\bf{E}}$ cannot necessarily be continued for all time, if $|{\bf{E}}(t)|$ approaches a positive multiple of $2\pi$ in finite time. If $|{\bf{E}}(t)|$ reaches say $23\pi/$ 2 at time $t_{1}$, restart with initial condition $-{\bf{E}}(t_{1})/3$, which amounts to flipping the axis and starting over with rotation angle $\pi/2$, corresponding to the same physical position. This avoids dealing with the boundary issues.
@@
$=1-\left({1-{{({\boldsymbol{\omega}}\cdot{{\bf{n}}_{0}})}^{2}}}\right){\sin^{2}}({\theta_{0}}/2)={a^{2}},$ so $\left|{{{\bf{v}}_{2}}}\right|=1,$ and so ${{\bf{v}}_{2}}={{\bf{n}}_{0}}({\boldsymbol{\omega}}\cdot{{\bf{n}}_{0}}/a)+{\bf{u}}\sqrt{1-{{({\boldsymbol{\omega}}\cdot{{\bf{n}}_{0}}/a)}^{2}}}.$ Similarly, ${{\bf{n}}_{0}}\sqrt{1-{{({\boldsymbol{\omega}}\cdot{{\bf{n}}_{0}})}^{2}}}\cos({\theta_{0}}/2)-{\bf{u}}({\boldsymbol{\omega}}\cdot{{\bf{n}}_{0}})$ has length $a$, so ${\left({1-{a^{2}}}\right)^{-1/2}}{{\bf{v}}_{1}}=\\ {{\bf{n}}_{0}}\sqrt{1-{{({\boldsymbol{\omega}}\cdot{{\bf{n}}_{0}}/a)}^{2}}}-{\bf{u}}({\boldsymbol{\omega}}\cdot{{\bf{n}}_{0}}/a).$ Thus ${\bf{n}}(t)\sin({\theta(t)}/2){\left({1-{a^{2}}}\right)^{-1/2}}$$={{\bf{e}}_{1}}\cos(t/2+b)+\\ {{\bf{e}}_{2}}{\left({1-{a^{2}}}\right)^{-1/2}}\sin(t/2+b)={\bf{p}}(t),$ and ${\bf{n}}(t)$ is the normalization of this.
Example.
Suppose the body starts out from having been rotated 90 degrees about the $x$ axis, and thereafter has angular velocity vector ${\bf{k}}$, so it is spinning about the vertical. That is, ${\theta_{0}}=\pi/2$, ${{\bf{n}}_{0}}={\bf{i}}$, and ${\boldsymbol{\omega}}={\bf{k}}$, and let ${t_{0}}=0$. Then in Theorem 4.1, $a=1/\sqrt{2}$ , $b=0$, and $\theta(t)=2{{\mathop{\rm Cos}\nolimits}^{-1}}\left({{2^{-1/2}}\cos\left({t/2}\right)}\right)$. Also, one can compute ${{\bf{e}}_{1}}={\bf{i}}$ and ${{\bf{e}}_{2}}=({\bf{j}}+{\bf{k}})/\sqrt{2}$, so ${\bf{n}}(t)=(\cos(t/2),\sin(t/2),\sin(t/2))/\sqrt{1+{{\sin}^{2}}(t/2)}.$ This is illustrated above in Figures 1 and 2. Note ${\bf{n}}$ starts at ${{\bf{i}}}$, picks up some positive 2nd and 3rd components, and then winds up at $-{{\bf{i}}}$ as $\theta$ goes from $\pi/2$ to $23\pi/$ 2, as the body makes one complete revolution. Then ${\bf{n}}$ picks up some negative 2nd and 3rd components, and winds up at ${{\bf{i}}}$ as $\theta$ goes from $23\pi/$ 2 back to $\pi/2$, as the body makes another complete revolution. Any example of Theorem 4.1 behaves similarly, except that the offset $b$ is not zero unless ${\boldsymbol{\omega}}$ is perpendicular to ${{\bf{n}}_{0}}$, so in general $\bf{n}$ does not start out at ${{\bf{e}}_{1}}$. See Figure 5.
The spinor-like behavior of Theorem 4.1, in which ${\bf{n}}$ changes sign when the body makes one complete revolution, does not actually require that ${\boldsymbol{\omega}}$ be constant in magnitude, only in direction. This is consequence of a simple change of time argument that is analogous to changing to arclength parameter for a curve.
Lemma 4.2.
@@
Thus in Theorem 4.1, the hypothesis could be changed to ${\boldsymbol{\omega}}/\left|{\boldsymbol{\omega}}\right|$ being constant, and the conclusions would hold using the new time variable, and the statement about ${\bf{n}}$ changing sign every time the body makes one complete rotation is still true. In general, the geometrical behavior of $\boldsymbol{E}$ is not affected by a change of time scale, in the same way that the geometry of a curve is not affected by a change of parameterization. This is useful in our computer experiments with $\boldsymbol{E}$: the magnitude of ${\boldsymbol{\omega}}$ might as well be kept fixed.
@@
Theorem 5.1 (Continuous evolution of Euler axis and angle).
@@
We’ll prove the theorem by analyzing the solution to equation (9), which exists for all time, at points where ${\bf{M}}={\bf{0}}$. Under the hypotheses of Theorem 5.1, these points are isolated. The trick will be to choose the correct branch of the $\arccos$ function in Theorem 3.2 between these points so that axis and angle connect up continuously.
Lemma 5.2 (Zeros are isolated).
Let ${\boldsymbol{\omega}}$ satisfy the hypotheses of Theorem 5.1, and let $({M_{0}(t),\bf{M}}(t))$ solve equation (9) for ${t_{0}}\leq t$, with $M_{0}(t_{0})^{2}+{\bf{M}}(t_{0})\cdot{\bf{M}}(t_{0})=1$. If $t_{0}<t_{1}$ and ${\bf{M}}(t_{1})={\bf{0}}$, there exists $\delta>0$ such that ${\bf{M}}(t)\neq{\bf{0}}$ for $0<|t-t_{1}|<\delta$.
@@
Lemma 5.3 (Continuation at a zero).
Let ${\boldsymbol{\omega}}$ satisfy the hypotheses of Theorem 5.1, and let
$({M_{0}(t),\bf{M}}(t))$ solve equation (9) for ${t_{0}}\leq t$, with $M_{0}(t_{0})^{2}+{\bf{M}}(t_{0})\cdot{\bf{M}}(t_{0})=1$. Suppose $t_{0}<t_{1}<t_{2}$, ${\bf{M}}(t_{1})={\bf{0}}$, and ${\bf{M}}(t)\neq{\bf{0}}$ if $t\neq t_{1}$ and $t_{0}<t<t_{2}$. Suppose, for some integer $l$, $\theta(t)$ has been defined for $t_{0}\leq t\leq t_{1}$ by $\cos(\theta(t)/2)=M_{0}(t)$ with $2\pi l\leq\theta(t)\leq 2\pi(l+1)$; and for $t_{0}<t<t_{1}$, ${\bf{n(t)}}={\bf{M}}(t)/\sin(\theta(t)/2)$. Let $i=min\{j\geq 0:{\boldsymbol{\omega}}^{(j)}(t_{1})\neq{\bf{0}}\}$.